help me solve f(x)=(1/4)^x Domain/Range/Horiz Asymptote
Answer by lwsshak3(11628)

help me solve f(x)=(1/4)^x Domain/Range/Horiz Asymptote
Start with a well-known point on all exponential functions, when x=0, y=1 (0,1)
When x>0, the denominator becomes larger and the fraction, f(x), approaches zero.
So the y-axis becomes a horizontal asymptote.
When x<0 , f(x)=(1/4)^-x=1/(1/4)^x=4^x, so f(x) increases to infinity.
Domain: (-∞,∞)
Range:(0,∞)
Asymptote: x-axis or y=0